OpenStudy (anonymous):

Simplify. –12 + 9 – (–3) – 11 A. –17 B. –11 C. 7 D. 13

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

i thought this was called "add" who wrote "simplify"?

OpenStudy (anonymous):

$-12+9-\left(-3\right)-11=-11$−12+9−(−3)−11=−11 Steps $-12+9-\left(-3\right)-11$−12+9−(−3)−11 $\mathrm{Follow\:the\:PEMDAS\:order\:of\:operations}$Follow the PEMDAS order of operations Show Steps $\mathrm{Calculate\:within\:parentheses}\:\left(-3\right)\::\quad-3$Calculate within parentheses (−3) : −3 $=-12+9-\left(-3\right)-11$=−12+9−(−3)−11 Show Steps $\mathrm{Add\:and\:subtract\:\left(left\:to\:right\right)}\:-12+9-\left(-3\right)-11\::\quad-11$Add and subtract (left to right) −12+9−(−3)−11 : −11 $=-11$=−11

OpenStudy (misty1212):

start with replacing \(-(-3)\) by \(+3\) and compute \[–12 + 9 +3 – 11\]

OpenStudy (anonymous):

hey where is the equation editor

OpenStudy (misty1212):

@usercode3rror wherever you are getting your answers from, it is not working you can't cut and paste with this code doe

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

$-12+9-\left(-3\right)-11=-11$−12+9−(−3)−11=−11

OpenStudy (anonymous):

=−12+9−(−3)−11

OpenStudy (anonymous):

=−11